This is an inroganic chemistry question, I just need help with number 7 but included everything so you can get background. Please only answer if you can answer it correctly.

a) Obtain two 50 mL beakers. Into one beaker add 1
Answer #1

QUESTION 7

When excess ammonia (or ammonium hydroxide) is added slowly into copper (II) nitrate, hydroxide ions are formed in the solution. As ammonia is a weak base, it removes hydrogen ions from water, to form hydroxide ions.

NH_3+H_2O\rightarrow NH_{4}^{+}+OH^-

First, these hydroxide ions will react to form an insolubale pale blue, copper (II) hydroxide.

Cu^{2+}+2OH^-\rightarrow Cu(OH)_2

Upon further addition, the colour of the solution turns dark blue, due to the formation of a copper-ammonia complex as follows.

Cu^{2+}+4NH_3\rightarrow [Cu(NH_3)_4]^{2+}

The drop in voltage can be attributed to the dilution of the cathodic electrolyte. When the copper solution is diluted with the ammonia solution, the concentration of the cathodic electrolyte decreases, and the voltage of the cell decreases. (The nernst equation can be used to explain this)

E_{cell}=E^0_{cell}-\frac{0.0592}{n}log\frac{[Zn^{2+}]}{[Cu^{2+}]}